# VISO: Robust Underwater Visual-Inertial-Sonar SLAM with Photometric Rendering for Dense 3D Reconstruction

### v1

@@ IV EXPERIMENTAL RESULTS
To validate the feasibility of the proposed VISO, we conducted extensive experimental evaluations in both a laboratory tank and an open lake. In all experiments, the state-of-the-art (SOTA) underwater SLAM algorithm SVIn2 [rahman2022svin2] and the visual–inertial odometry algorithm VINS-Fusion [qin2017vins] are used as baselines for comparison. Since a profiling sonar is prohibitively expensive and not available in our setup, only the camera and IMU are used, hence denoted as SVIn2 (VI). Apart from the SOTA visual SLAM algorithms, we also compare against the Dead Reckoning approach [moore2015generalized], which fuses measurements from an Attitude and Heading Reference System (AHRS), a depthometer, a compass, and a DVL using an Extended Kalman Filter (EKF).
@@ IV-A Lab Tank Experiments
We first conducted experiments in a laboratory tank measuring $12\times 12m$ with a depth of $2.85m$. A Qualisys underwater motion capture system, equipped with four cameras mounted around the tank, was used to provide ground-truth, as shown in Fig. 5(b). The robot platform is a Bluerov2 with an eight-thruster configuration. It is equipped with an AHRS, a customised stereo camera, a WaterLinked Sonar 3D-15 imaging sonar, a Nortek Nucleus 1000 DVL that contains an IMU, a depth sensor, and an altitude sensor, as shown in Fig. 5(a). In all experiments, we use the IMU located in the Nortek Nucleus 1000 DVL instead of the onboard IMU of the Bluerov2.
@@ IV-A3 Dense Mapping Experiment
In this work, we propose a real-time novel dense mapping approach for underwater 3D scene reconstruction using 3D sonar data and photometric rendering. To evaluate its performance, we conducted a mapping experiment in the laboratory tank and compared our approach with COLMAP [schonberger2016structure], a widely used structure-from-motion (SFM) method for offline dense mapping. As shown in Fig. 8, our method, which exploits rendered 3D sonar point clouds for dense mapping, achieves performance comparable to the SOTA offline algorithm. However, while SFM requires approximately 20 minutes on a server equipped with 8 TITAN X GPUs and a 32-core Intel(R) Xeon(R) CPU, our method generates the map in real time.

### v2

@@ IV EXPERIMENTAL RESULTS
To validate the feasibility of the proposed VISO, we conducted extensive experimental evaluations in both a laboratory tank and an open lake. In all experiments, the state-of-the-art (SOTA) underwater SLAM algorithm SVIn2 [20] and the visual–inertial odometry algorithm VINS-Fusion [18] are used as baselines for comparison. Since a profiling sonar is prohibitively expensive and not available in our setup, only the camera and IMU are used, hence denoted as SVIn2 (VI). Apart from the SOTA visual SLAM algorithms, we also compare against the Dead Reckoning approach [15], which fuses measurements from an Attitude and Heading Reference System (AHRS), a depthometer, a compass, and a DVL using an Extended Kalman Filter (EKF).
@@ IV-A Lab Tank Experiments
We first conducted experiments in a laboratory tank measuring $12\times 12m$ with a depth of $2.85m$. A Qualisys underwater motion capture system, equipped with four cameras mounted around the tank, was used to provide ground-truth, as shown in Fig. 5(b). The robot platform is a Bluerov2 with an eight-thruster configuration. It is equipped with an AHRS, a stereo camera built by Frontier Robotics, a WaterLinked Sonar 3D-15 imaging sonar, a Nortek Nucleus 1000 DVL that contains an IMU, a depth sensor, and an altitude sensor, as shown in Fig. 5(a). In all experiments, we use the IMU located in the Nortek Nucleus 1000 DVL instead of the onboard IMU of the Bluerov2.
@@ IV-A3 Dense Mapping Experiment
In this work, we propose a real-time novel dense mapping approach for underwater 3D scene reconstruction using 3D sonar data and photometric rendering. To evaluate its performance, we conducted a mapping experiment in the laboratory tank and compared our approach with COLMAP [21], a widely used structure-from-motion (SFM) method for offline dense mapping. As shown in Fig. 8, our method, which exploits rendered 3D sonar point clouds for dense mapping, achieves performance comparable to the SOTA offline algorithm. However, while SFM requires approximately 20 minutes on a server equipped with 8 TITAN X GPUs and a 32-core Intel(R) Xeon(R) CPU, our method generates the map in real time.